From a large number of actuarial exam scores, a random sample of 250 scores is selected, and it is found that 177 of these 250 are passing scores. Based on this sample, find a 90% confidence interval for the proportion of all scores that are passing. Then complete the table below.
Carry your intermediate computations to at least three decimal places. Round your answers to two decimal places.
What is the lower limit of the 90% confidence interval?
What is the upper limit of the 90% confidence interval?

Answer by stanbon(75887)

From a large number of actuarial exam scores, a random sample of 250 scores is selected, and it is found that 177 of these 250 are passing scores.
--------------------------
Based on this sample, find a 90% confidence interval for the proportion of all scores that are passing. Then complete the table below.
--------------------------
p-hat = 177/250 = 0.708
ME = 1.645*sqrt[0.708*0.292/250] = 0.047
------------------

Carry your intermediate computations to at least three decimal places. Round your answers to two decimal places.
What is the lower limit of the 90% confidence interval?::0.71-0.05 = 0.66
What is the upper limit of the 90% confidence interval?::0.71+0.05 = 0.76
============================
Cheers,
Stan H.
=============